OpenStudy (anonymous):

how do i solve the equation [x^2 + (2x^2)y - 4y^3 = xy]. using y' = dy/dx? urgent i got a test on this in 4 hours

OpenStudy (anonymous):

by solve what do you mean? there are lots of pairs of numbers (x,y) satisfying that, do you mean differentiate?

OpenStudy (anonymous):

do you want to fing y' ?

OpenStudy (anonymous):

find

OpenStudy (anonymous):

what do you want us to do

OpenStudy (anonymous):

the question was to find y' = dy/dx of the equation by implicit diffrentation

OpenStudy (anonymous):

differentiate on both sides use product rule where ever necessary the answer would be \[2x+4xy+2x ^{2}y \prime-12y ^{2}y \prime=xy \prime+y\] now solve this for \[y \prime=(y-2x-4xy)/(2x ^{2}-12y ^{2}-x)\]

OpenStudy (experimentx):

then find y' using implicit differentiation ... http://www.youtube.com/watch?v=sL6MC-lKOrw

OpenStudy (anonymous):

?

OpenStudy (anonymous):

explain the final parts please..

OpenStudy (experimentx):

??

OpenStudy (anonymous):

dont get it

OpenStudy (experimentx):

in the video or in the above problem??

OpenStudy (anonymous):

the problem

OpenStudy (experimentx):

fist step .. differentiate second step ... take y' common <--- separate y' on one side!!

OpenStudy (anonymous):

(y* dy/dx - 2 - 4y*dy/dx) / (4x-24*dy/dx-1) ??

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

first off we will differentiate: \[x^2 + 2x^2y - 4y^3 = xy\] the first term: \[\frac{d}{dx} (x^2) = 2x\] the second term uses the product rule: \[\frac{d}{dx} ( 2x^2 y) = 4xy + 2x^2 \frac{dy}{dx}\] the third term uses the chain rule: \[\frac{d}{dx} ( -4y^3) = -12y^2 \frac{dy}{dx}\] on the other side we have another product rule: \[\frac{d}{dx} ( xy )= y + x \frac{dy}{dx}\] so differentiating the equation \[x^2 + 2x^2y - 4y^3 = xy\] we get \[2x + 4xy +2x^2 \frac{dy}{dx} - 12x^2 \frac{dy}{dx} = y + x \frac{dy}{dx}\]

OpenStudy (anonymous):

are you ok with that?

OpenStudy (anonymous):

i'll continue, if anything is confusing just ask and i'll explain more :)

OpenStudy (anonymous):

we now want to isolate dy/dx so just as if we were isolating anything else, we want to collect "like" terms, factorising and stuff: \[2x + 4xy +2x^2 \frac{dy}{dx} - 12x^2 \frac{dy}{dx} = y + x \frac{dy}{dx}\] \[2x^2 \frac{dy}{dx} - 12x^2 \frac{dy}{dx} - x \frac{dy}{dx} = y -2x - 4xy\]\[\frac{dy}{dx} (2x^2 - 12x^2 - x ) =( y -2x - 4xy)\]\[\frac{dy}{dx} =\frac{ y -2x - 4xy}{2x^2 - 12x^2 - x }\]

OpenStudy (anonymous):

warning: mistake!

OpenStudy (anonymous):

that -12x^2 should be a -12y^2

OpenStudy (anonymous):

\[\frac{dy}{dx} =\frac{ y -2x - 4xy}{2x^2 - 12y^2 - x }\]

OpenStudy (anonymous):

hate typing